OpenStudy (anonymous):

Albert could race 4 times as fast as Bernard. For this reason Albert could race 48 miles in 3 hours less than it took Bernhard to race 3o miles. How fast could each of them race? What was the time of each racer?

OpenStudy (anonymous):

Let A and B be the speed of Albert and Bernard respectively. A=4B t = d/r \[\frac{48}{A}=\frac{30}{B}-3 \] Solve the two simultaneous equations for A and B \[\{A=24,B=6\} \]Can you complete the problem from here?

OpenStudy (anonymous):

No, I don't know how...

OpenStudy (anonymous):

Sorry, I have to sign off.